OpenStudy (anonymous):

A baby whale was growing at an average rate of 250 pounds per day. What is the equivalent of this rate in ounces per hour?

OpenStudy (anonymous):

16 ounces=1pound 1 day=24 hours

OpenStudy (anonymous):

250 divided by 24= pounds per hour then you take the answer from that and divide it by 16 and you will know how many ounces it grows every hour.

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

Write each conversion factor as a fraction that equals 1. Then multiply the given number by the conversion factors that will cancel out the unwanted units and will keep the desired units.

OpenStudy (mathstudent55):

Here is an example. What is 50 miles per hour in feet per second? 1 mile = 5280 ft \(\dfrac{1~mile}{5280~ft} = \dfrac{5280~ft}{1~mile} = 1\) 1 hour = 3600 seconds \(\dfrac{1~hour}{3600~seconds} = \dfrac{3600~seconds}{1~hour} = 1\)

OpenStudy (mathstudent55):

Now that we have the conversion factors, we use the appropriate ones: \(50 \dfrac{\cancel{miles}}{\cancel{hour}} \times \dfrac{5280~ft}{1~\cancel{mile}} \times \dfrac{1~\cancel{hour}}{3600~seconds} \) Notice that the only units left are ft/sec. The answer is \(73 \dfrac{1}{3} \dfrac{feet}{sec} \)

OpenStudy (anonymous):

250 pounds per day=250*16=4000 ounces per day=4000/24=500/3=166 1/3 ounces per hour

OpenStudy (mathstudent55):

\(250\dfrac{lb}{day} \times \dfrac{16~oz}{lb} \times \dfrac{1~day}{24~hours} \) \(=250\dfrac{\color{red}{\cancel{lb}}}{\color{green}{\cancel{day}}} \times \dfrac{16~oz}{\color{red}{\cancel{lb}}} \times \dfrac{1~\color{green}{\cancel{day}}}{24~hours} \) \(= 166 \dfrac{2}{3} \dfrac{oz}{hour}\)